OpenStudy (anonymous):

Please help me! will medal and fan! thank you♥ The diameter of a tire is 2.5 ft. Use this measurement to answer parts a and b. Show all work to receive full credit. a. Find the circumference of the tire. Round your answer to the nearest whole number. b. About how many times will the tire have to rotate to travel 1 mile? (1 mile = 5,280 ft)

OpenStudy (vijeya3):

diameter=2.5ft so,radius=2.5/2 =25/20 now the circumference is given be 2*pi*r therefore circumference=0.31425 ft(approx) nearest whole number should be 1(since it cannot be 0!) This was the A part For B part... let the number of rotations be x so x*1ft=5280ft x=5280 so the tire rotates 5280 times.... If this is helpful,fan and medal me..:)